OpenStudy (anonymous):

Fencing is sold in 8 foot sections at $15 per section. Fence posts cost $4 each. How much does it cost to build a fence around a rectangular field which is 240 feet wide and 320 feet long?

OpenStudy (anonymous):

find the perimeter: P=240*2+320*2 P=1120 divide that by 8 sections: 1120/8=140 multiply by $15 per section: 140*15= 2100 find how many posts you need: 30+30+40+40=140 multiply number of posts by $4 a post: 140 *4=560 add everything up: 2100+560 Grand Total of $2660